# Evaluating Sample Utility for Data Selection by Mimicking Model Weights

### v2

@@ Abstract
Foundation models are trained on large-scale web-crawled datasets, which often contain noise, biases, and irrelevant information. This motivates the use of data selection techniques, which can be divided into model-free variants—relying on heuristic rules and downstream datasets—and model-based, e.g., using influence functions. The former can be expensive to design and risk introducing unwanted dependencies, while the latter are often computationally prohibitive. Instead, we propose an efficient, model-based approach using the Mimic Score, a new data quality metric that leverages the weights of a reference model to assess the usefulness of individual samples for training a new model. It relies on the alignment between gradients and a target direction induced by the reference model. Using the derived Mimic Scores, we develop Grad-Mimic, a framework that prioritizes samples for learning, creates effective filters, and automates data selection. Empirically, using Mimic Scores to guide training improves data efficiency, results in consistent performance gains across six image datasets, and includes enhancements to CLIP models. Moreover, Mimic Score-based filters improve upon existing filtering methods, e.g., cutting 4.7 million samples to train better CLIP models while offering accurate estimation of training dataset quality.
@@ 1 Introduction
Large-scale web-crawled datasets are fundamental to the success of foundation models, e.g., MM1.5 [1], OpenAI CLIP [2], and GPT-4 [3]. These datasets provide vast quantities of information but also carry noise, biases, and irrelevant content. To mitigate these, data selection—ruling out undesirable samples—has emerged as a critical step in the development pipeline [4, 5]. For example, the FineWeb dataset [6], with 15 trillion tokens, is refined through eight carefully designed filtering steps to ensure high model performance.
Existing data selection methods have several limitations. We can divide these into two categories: those that are independent of a model being trained (or using pretrained models) and those that rely on such a model. Model-free approaches, such as those based on handcrafted heuristics or semantic similarity to downstream datasets [7, 8] often require expensive experimentation to establish selection rules and add unwanted dependencies on additional datasets. Furthermore, these methods can fail to provide fine-grained insights into individual samples and overlook sample utility in training, resulting in suboptimal performance. In contrast, model-based techniques, e.g., training specialized filtering networks  [9, 10, 11, 12] or using data influence functions  [13, 14], introduce complexity and are often computationally expensive.
@@
Figure 1: High-value vs. low-value samples identified by the Mimic score: We present randomly selected samples from the top 5% (first row) and bottom 5% (second row) of web-crawled data, ranked by their Mimic scores. Each image is accompanied by its caption and corresponding CLIP score (higher scores suggest better quality). Mimic scores align closely with CLIP scores. High-value samples generally have detailed captions and coherent visual content, while low-value ones carry short captions and misaligned content.
To address these limitations, we introduce Mimic Score, a new data quality metric that assesses sample’s contribution to weight updates using reference model’s weights. This is an efficient model-based technique. We show that by leveraging pretrained model weights (located in a more optimal part of the weight space), we bypass the need for downstream datasets or specialized model training and can use it as our selection guide. Our approach is based on the alignment between each sample’s negative gradient and a direction pointing toward the reference model in the weight space.
We argue that samples that potentially pull the model in undesirable directions—thereby misguiding weight updates—should be considered low-value and ruled out.
Leveraging model weights to evaluate sample utility offers several advantages. First, with the growing availability of well-trained models, adopting model weights has become more accessible than ever. Second, unlike other reference model-based approaches that require loading a second model and performing full inference passes [15, 16], our method scores each sample directly from the weights, improving efficiency.
@@
Building on the mimic score, we develop Grad-Mimic, a data selection framework that operates in two stages:
  • Stage 1: During training, Grad-Mimic prioritizes which samples to learn based on their mimic scores, improving data efficiency while steering the model toward the reference model’s weight space.
  • Stage 2: After training, Grad-Mimic uses mimic scores to evaluate each sample’s utility across training steps. These assessments are then aggregated to create an ensemble filter, which can further refine the dataset.
@@
We validate Grad-Mimic through diverse experiments. First, in a controlled setting with varying levels of label noise, Grad-Mimic accurately identifies mislabeled samples and estimates training dataset quality. During training, it effectively down-weights undesirable samples and improves data efficiency across six image datasets. Next, we test it on large-scale web-crawled multimodal data by training CLIP models from scratch [2] on 10M and 100M samples from DataComp  [17]. Using publicly available pretrained weights as our reference, mimic scores effectively navigate training. The resulting filter performs better than human-designed ones and can complement CLIP score-based filters [18], lifting CLIP model performance with less training data.
@@
  • New Data Quality Metric: We propose a novel approach to assess sample utility by computing the alignment between gradients and the vector toward the reference model in the weight space.
  • Automated Data Selection Framework: Using the mimic score, we present Grad-Mimic, a framework that effectively identifizes useful samples to improve data efficiency and automates effective data selection.
  • Dataset Quality Assessment: The mimic score, along with its resulting filters, improves existing filtering strategies and provides reliable estimation of dataset quality.
@@ 2 Related Work
Figure 2: Grad-Mimic Two-stage Workflow: In Stage 1, Grad-Mimic uses the vector pointing toward the reference model’s weight space to measure alignments with each sample’s negative gradients. Then it adapts alignment scores to reweight gradients that contributed to the weight updates. In Stage 2, derived alignment scores—mimic scores—are used to identify low-value samples and build an ensemble filter, which can be used for noise detection, dataset quality estimation, and training dataset discovery.
Data Selection.
Data selection techniques can be categorized into group- and sample-level approaches. Group-level approaches focus on optimizing data domain mixtures to curate high-quality datasets [19, 20, 21]. Sample-level methods, which are the focus of Grad-Mimic, aim to provide a score measuring sample utility and filter out unhelpful samples. Prior research has explored various gradient-based strategies, such as identifying impactful samples through gradient magnitudes [22], analyzing gradient similarities across batches [23], and selecting key samples that can capture full training [24, 25]. These methods can struggle with large amounts of noise in batch gradients. Grad-Mimic overcome this with the help of a reference model.
Data Curation for Multimodal Models.
Multimodal models rely on large-scale web datasets [26, 17, 27], which are often naturally noisy and require careful curation. Previous approaches have included selecting samples based on semantic similarity to the downstream evaluation datasets [7, 8], using semantic deduplication [28], developing specialized filtering networks [9, 10, 11, 12], and using data influence functions [13, 14]. While effective, they require access to target datasets, introduce training complexities, and can be expensive in computation. Grad-Mimic overcomes these and offers a more efficient alternative: using public available model weights to identify useful samples.
Weak Supervision.
Weak supervision is an emerging paradigm for constructing labeled datasets by combining multiple noisy label estimates [29, 30, 31, 32]. These estimates can come from sources such as heuristic labeling rules, domain knowledge, or pretrained models [33, 34], often encoded as labeling functions. The outputs of these labeling functions are modeled and then aggregated to produce final labels [29]. Weak supervision has demonstrated success across diverse domains [35, 36, 37, 38, 39]. Most prior works focus on weak label aggregation to construct datasets, Grad-Mimic leverages weak supervision for a new purpose: modeling sample utility assessments for a combined filtering decision.
@@ 3 Evaluating Sample Utility
We quantify samples by their contributions to the learning process. Our principle is that samples that potentially pull the model in an undesirable direction, thereby misdirecting weight updates, should be considered low-value.
We start with setup and notation, then explain how the scoring metric is derived and used in our data selection framework, Grad-Mimic.
Notation.
Let $D=\{s_{i}\}_{i=1}^{n}$ denote a dataset of $n$ samples drawn from a distribution $\mathcal{D}$ supported on the space $\mathcal{S}$. At training step $t$, model parameters $\theta_{t}$, are iteratively optimized using the dataset $D$. While our framework supports various training settings, we focus on supervised learning for clarity. We assume $\mathcal{S}=\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ is the input space and $\mathcal{Y}$ is the label space. Each sample $s_{i}$ can be expressed as $(x_{i},y_{i})$, where noise may be present either in the instance $x_{i}$ or in the label $y_{i}$. The empirical loss across samples is defined as $\frac{1}{n}\sum_{i=1}^{n}\ell(x_{i},y_{i})$, and each sample’s gradient with respect to the model parameters $\theta_{t}$ is written as: $g_{i,t}:=\nabla_{\theta_{t}}\ell(x_{i},y_{i})$. A standard update to model parameters is $\theta_{t+1}:=\theta_{t}-\eta\sum_{i=1}^{b}g_{i,t}$, where $\eta$ is the learning rate and $b$ is the batch size.
3.1 Mimic Score Calculation
To evaluate whether a sample’s gradient pulls the model in an undesirable direction, we use the weights of a pretrained reference model as our guide. These reference weights, denoted by $\theta_{\text{ref}}$, reside in a more optimal part of the weight space, such that $\ell(\theta_{\text{ref}})<\ell(\theta_{t})$. We use the vector from the current model’s weight space $\theta_{t}$ to $\theta_{\text{ref}}$ to measure each sample utility in approximating a better weight configuration ($\theta_{\text{ref}}$). These weights can be layer-specific, e.g., model weights in the last layer, which usually store more informative features [40, 41].
Reference weights can be obtained in two ways: by training a model on the dataset to achieve a desired performance level, if resources permit, or more efficiently, by using publicly available well-trained models, which are now more accessible and can eliminate the need for direct data access or possible costly training.
@@ 3.1 Mimic Score Calculation
This alignment score, named the mimic score, reflects how much a sample can drive the model closer to the reference model. A sample having a lower mimic score suggests it has limited utility in guiding model updates, making it a potential candidate for exclusion from future training.
4 Grad-Mimic: Data Selection Framework
Building on this simple but effective metric, we propose Grad-Mimic, a two-stage data selection framework that first prioritizes which samples to learn and then identifies useful samples to design an effective filter. The complete Grad-Mimic workflow is illustrated in Fig. 2.
4.1 Stage 1: Reweighted Gradient Updates
We first use mimic scores to aid model training by re-weighting sample gradients. Unlike standard gradient descent, which assigns equal weight to all the samples in the mini-batch, Grad-Mimic uses mimic scores to amplify helpful samples and down-weight unhelpful ones.
@@ 4.1 Stage 1: Reweighted Gradient Updates
The temperature $\tau$ controls the sensitivity of sample reweighting, allowing us to adjust how sharply the model prioritizes samples. A lower temperature results in a more aggressive focus on learning the most aligned samples, while a higher temperature encourages Grad-Mimic converges to standard gradient descent.
4.2 Stage 2: Automated Filter Design
The second stage of Grad-Mimic uses computed mimic scores to identify valuable samples, creating a combined filter based on sample utility assessments.
4.2.1 Sample Identification
Grad-Mimic first gathers normalized mimic scores for each sample at every training step. These scores indicate the sample contributions over training, which allows us to decide whether to retain or discard a sample. Grad-Mimic supports several sample identification methods:
  • Threshold-based: A sample is chosen if its mimic score exceeds a defined threshold. This threshold could be set as $1/b$ (indicating greater than uniform weight).
  • 1D Clustering: Samples are categorized into two groups using clustering such as $k$-means clustering [42] or Gaussian Mixture Models (GMMs), allowing more automated identification based on their assigned cluster.
@@ 4.2.1 Sample Identification
After identification, normalized mimic scores are binarized, and samples are categorized into “retain” or “discard” at each training step. We treat these assigned groups as each training step’s votes and then combine for a final filter.
4.2.2 Filter Output Aggregation
Although using the assessment at a training step can refine the dataset, relying on such isolated filters risks overlooking training dynamics. Moreover, evaluating a sample’s utility across all training stages may yield complementary signals.
To address this, Grad-Mimic uses weak supervision techniques [29, 30, 31, 32], which are usually used for constructing datasets from noisy signals, to combine filter outputs across training steps. We start by learning a model to evaluate the reliability of each step’s assessment [31]. Once established, this model aggregates filtering decisions into a high-quality ensemble filter. Through aggregation, Grad-Mimic mitigates assessment noise and captures training dynamics. We employ the Snorkel framework [30], a widely used method in the weak supervision community.
Ultimately, we can use this ensemble filter to select high-value samples for a refined dataset. Moreover, we can use it to estimate training dataset quality by either analyzing the proportion of retained samples or by averaging samples’ mimic scores as an overall metric. We summarize the steps of Grad-Mimic in Algorithm 1 and detailed in Appendix A.

### v3

# Evaluating Sample Utility for Efficient Data Selection by Mimicking Model Weights

@@ Abstract
Multimodal models are trained on large-scale web-crawled datasets, which often contain noise, bias, and irrelevant information. This motivates the use of data selection techniques, which can be divided into model-free variants, relying on heuristic rules and downstream datasets, and model-based approaches, such as those using influence functions. The former can be expensive to design and risks introducing unwanted dataset dependencies, while the latter are often computationally prohibitive. In this work, we propose an efficient, model-based approach using the Mimic Score, a new data-quality metric that leverages the weights of a reference model to assess the usefulness of individual samples for training a new model. Our method relies on measuring alignments between training gradients and a target direction induced by this reference model. Building on the derived mimic scores, we develop Grad-Mimic: a framework that prioritizes samples to learn, estimates overall sample utility, and creates effective filters. Empirically, using mimic scores to guide training improves data efficiency, accelerates convergence, yields consistent performance gains across six image datasets, and enhances CLIP models with 20.7% fewer training steps. Moreover, mimic score-based filters complement existing filtering methods, e.g., training improved CLIP models with 4.7 million fewer samples while offering accurate estimation of dataset quality.
@@ 1 Introduction
Large-scale web-crawled datasets are fundamental to the success of modern multimodal models such as AIMv2 [1], ALIGN [2], OpenAI CLIP [3], and SigLIP [4]. These datasets provide vast quantities of information—but inevitably contain noise, biases, and irrelevant content. To mitigate these, data selection—ruling out undesirable samples—has emerged as a critical step in the development pipeline [5, 6]. For example, the FineWeb dataset [7], containing 15 trillion text tokens, undergoes eight carefully designed filtering steps to ensure high model performance.
Current selection strategies; however, are far from satisfactory. We can broadly divide them into two categories: model-free methods, such as those based on handcrafted heuristics or semantic similarity to downstream datasets [8, 9] often require costly trial-and-error to establish selection rules and add unwanted dependencies on other datasets. These methods also lack the granularity needed to assess individual utility, leading to suboptimal training outcomes. In contrast, model-based techniques, such as those building specialized filtering networks [10, 11, 12, 13], scoring samples through a pretrained reference model’s loss [14, 15], or using influence functions  [16, 17, 18, 19, 20], increase the complexity of the pipeline and are often computationally expensive. These shortcomings motivate the need for a simpler, dataset-agnostic, compute-efficient way to curate high-quality datasets.
@@
Figure 1: High-value vs. low-value samples identified by the Mimic score: Randomly selected samples from the top 5% (first row) and bottom 5% (second row) of web-crawled data, ranked by their mimic scores. Each image includes its caption and CLIP score (higher indicates better quality). Mimic scores align closely with CLIP scores. High-value samples generally have detailed captions and coherent visuals, while low-value ones carry short captions and misaligned content.
We introduce the Mimic Score, a new data-quality metric that quantifies each training sample’s contribution to effective weight updates. We argue that samples whose gradients potentially pull the model in undesirable directions—and thus misguide weight updates—should be down-weighted or discarded. To establish a desirable weight update, we leverage the direction pointing toward the weights of a pretrained reference model111 Note that our goal is to select high-quality data, not to reproduce the reference model. While insight into reference model’s training data would be valuable, we assume it is inaccessible—a common situation with many modern open-source models. , located in a more optimal part of the weight space, as our proxy. The mimic score is computed for each sample by measuring the alignment between its negative gradient and the direction induced by this reference model. A high alignment score indicates a sample can steer the model toward this preferred region, while a low one would divert the model.
Using model weights to evaluate sample utility offers multiple practical benefits. First, the growing supply of well-trained models makes this approach more accessible than ever. Second, unlike methods that derive selection guides using the gradients on noise-free validation datasets [18, 19, 20], mimic score avoids extra validation computations, bypassing the difficulties of curating such dataset and ensuring its quality. Finally, many reference model-based techniques compute excess loss through additional full inference [14, 15]. Our mechanism instead is a much simpler operation based on the discrepancy of model weights, which can substantially reduce compute overheads.
@@
Building on the mimic score, we develop Grad-Mimic, a two-stage data-selection framework:
  • Online Batch Re-weighting. During training, Grad-Mimic computes mimic scores on the fly and prioritizes samples to learn, enhancing data efficiency and accelerating convergence.
  • Offline Sample Selection. After training, Grad-Mimic aggregates sample’s mimic scores across training steps to estimate an overall sample utility. Such combined estimate serves as a reliable filter, enabling the construction of a smaller, higher-quality dataset for future use.
@@
We validate the capabilities of Grad-mimic in a broad range of scenarios. For mislabeled sample detection, we show in a controlled setting, Grad-Mimic accurately identifies noisy labels and estimates overall dataset quality. By down-weighting undesirable samples during training, it enhances data efficiency and model performance across six image datasets. For large-scale data curation, we scale our testbed to web-crawled multimodal data by training CLIP models from scratch [3] on 10M and 100M samples [21]. Using publicly released pretrained weights as our reference, mimic scores effectively navigate training, reducing 20.7% training steps to convergence. The resulting filter outperforms human-designed ones and complements CLIP score-based filters [22], lifting CLIP model performance with 4.7M fewer training samples.
@@
  • New Data-quality Metric: We introduce Mimic Score to assess sample utility by measuring the alignment between training gradients and a vector toward the reference model in the weight space.
  • Efficient Data Selection Framework: Built on Mimic Score, we present Grad-Mimic, a framework that efficiently prioritizes valuable samples to learn, enhances data efficiency, accelerates convergence, and automates effective data selection.
  • Dataset Quality Assessment: Mimic Score, along with its ensemble filters, provides reliable estimates of dataset quality and model performance gains and improves existing filtering strategies.
@@ 2 Related Work
Our work explores using the weights of a reference model to evaluate sample utility, intersecting three key areas: (i) data selection, (ii) data curation for multimodal models, and (iii) weak supervision.
Figure 2: Grad-Mimic Two-stage Workflow: During training, Grad-Mimic measures alignments using the vector toward the reference model’s weight space against each sample’s negative gradient. It then adapts alignment scores to re-weight gradients for weight updates. After training, the derived alignment scores, mimic scores, are used to identify low-value samples and build an ensemble filter, which can be used for noise detection, dataset quality assessment, and training dataset discovery.
Data Selection. Data selection techniques can be categorized into group- and sample-level approaches. Group-level approaches focus on optimizing data domain mixtures to curate high-quality datasets [23, 24, 25, 26]. Sample-level methods, which are the focus of Grad-Mimic, aim to provide a score measuring sample utility and rule out undesired samples. Prior research has explored various gradient-based strategies, such as identifying impactful samples through gradient magnitudes [27], analyzing gradient similarities to other training batches [28], selecting key samples that can capture full training [29, 30], and using gradients computed on noisy-free validation sets [18, 19, 20]. These methods; however, are sensitive to noisy batch gradients and often incur extra computation. Grad-Mimic sidesteps both issues with the help of a reference model.
Data Curation for Multimodal Models. Multimodal models rely on large-scale web datasets [31, 21, 32]; however, quantity comes at a price: unconstrained crawling amplifies noise and requires careful curation. Previous approaches have included selecting samples based on semantic similarity to the downstream evaluation datasets [8, 9], using semantic deduplication [33], developing specialized filtering networks [10, 11, 12, 13], and using data influence functions [16, 17]. While effective, they require access to other datasets, introduce training complexities, or demand substantial compute. Grad-Mimic offers a more efficient alternative: using public available models and the discrepancy in model weights to identify useful samples.
Weak Supervision. Weak supervision is a framework to construct labeled datasets by combining multiple noisy label estimates [34, 35, 36, 37]. These estimates can come from sources such as heuristic labeling rules, domain knowledge, or pretrained models [38, 39], often encoded as labeling functions. Labeling functions’ outputs are modeled and aggregated to produce probabilistic labeling decisions [34]. Weak supervision has demonstrated success in diverse domains [40, 41, 42, 43, 44]. Most prior works focus on weak label aggregation to construct datasets. Grad-Mimic leverages it for a new purpose: modeling each sample’s utility assessments for a combined filtering decision.
@@ 3 Evaluating Sample Utility
Our goal is to quantify training sample contributions to learning process when we (i) have a model to be trained and (ii) access a pretrained reference model as proxy. Our principle is that samples that potentially pull the model in an undesirable direction, thereby misdirecting weight updates, should be considered low-value. We start with setup and notation, then explain how an intermediate scoring metric is derived and used inside our data selection framework, Grad-Mimic (Sec. 4).
Notation. Let $D=\{s_{i}\}_{i=1}^{n}$ denote a dataset of $n$ samples drawn from a distribution $\mathcal{D}$ supported on the space $\mathcal{S}$. At training step $t$, model parameters $\theta_{t}$, are iteratively optimized using the dataset $D$. While our framework supports various training settings, we focus on supervised learning for clarity. We assume $\mathcal{S}=\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ is the input space and $\mathcal{Y}$ is the label space. Each sample $s_{i}$ can be expressed as $(x_{i},y_{i})$, where noise may be present either in the instance $x_{i}$ or in the label $y_{i}$. The empirical loss across all the samples is defined as $\frac{1}{n}\sum_{i=1}^{n}\ell(x_{i},y_{i})$, and each sample’s gradient with respect to the model parameters $\theta_{t}$ is written as: $g_{i,t}:=\nabla_{\theta_{t}}\ell(x_{i},y_{i})$. A standard update to model parameters is $\theta_{t+1}:=\theta_{t}-\eta\frac{1}{b}\sum_{i=1}^{b}g_{i,t}$, where $\eta$ is the learning rate and $b$ is the batch size.
Mimic Score Calculation. To evaluate whether a sample’s gradient pulls the model in an undesirable direction, we use the weights of a pretrained reference model as our proxy guide. These reference weights, denoted by $\theta_{\text{ref}}$, are assumed reside in a more optimal part of the weight space, such that $\ell(\theta_{\text{ref}})<\ell(\theta_{t})$. We use the vector from the current model’s weight space $\theta_{t}$ to $\theta_{\text{ref}}$ to measure each sample utility in approximating a better weight configuration ($\theta_{\text{ref}}$). These weights can be layer-specific, e.g., model weights in the last layer, which usually store more informative features [45, 46].
Reference weights can be obtained either by training a model on a (hold-out) dataset until it meets the desired performance [15, 14], if resource permits, or, more efficiently, by using an existing well-trained public model, which sidesteps both data-access constraints and costly training.
@@
This alignment score, named mimic score, reflects how much a sample can drive the model closer to a preferred weight space. A sample having a lower mimic score suggests it has limited utility in guiding effective weight updates, making it a potential candidate for exclusion from future training.
4 Data Selection Framework
Building on this simple data-quality metric, we package it into Grad-Mimic, a two-stage data selection framework that first prioritizes which samples to learn (Sec. 4.1) and then combines sample utilities to design an effective ensemble data filter (Sec. 4.2). A complete Grad-Mimic workflow is in Fig. 2.
4.1 Stage 1: Online Batch Re-weighting.
We first use mimic scores to aid model training through improving data efficiency by re-weighting. Unlike gradient descent, which assigns uniform weight to mini-batch samples, Grad-Mimic computes mimic scores on the fly to amplify helpful samples and down-weight unhelpful ones.
@@ 4 Data Selection Framework
The temperature $\tau$ controls the sensitivity of sample reweighting, allowing us to adjust how sharply the model prioritizes samples. A lower temperature results in a more aggressive focus on learning the most aligned samples, while a higher one encourages Grad-Mimic converges to gradient descent.
Theoretical Analysis. We provide an analysis for effective batch reweighting in Appendix A. In particular, we study conditions where Grad-Mimic converges faster than standard gradient descent.
4.2 Stage 2: Offline Sample Selection.
The second stage of Grad-Mimic converts computed mimic scores, estimated sample utilities, into a unified assessment, enabling us to curate a higher-quality training dataset with developed data filter.
Mimic Score Binarization. Grad-Mimic first gathers normalized mimic scores for each sample at every training step. These indicate their intermediate contributions to learning. We convert continuous scores into binary retain/discard decisions and explore three schemes:
  • Threshold-based: A sample is chosen if its mimic score exceeds a defined threshold. A natural choice can be set as $1/b$ (indicating greater than uniform).
  • 1D Clustering: Samples are categorized into two groups using clustering (e.g., $k$-means [47] or Gaussian mixture), allowing unsupervised identification based on their assigned clusters.
@@
Decision Aggregation. A filtering decision made at a single step can be noisy and ignore training dynamics. To address this, we treat each step’s binarized assessments as weak votes and use weak supervision techniques [34, 35, 36, 37] to combine filtering decisions across training steps. We start by learning a generative model to estimate the reliability of each step’s assessments [36]. Once established, this model produces probabilistic aggregated decisions. This consensus mitigates assessment noise and captures how a sample’s utility evolves. Empirically, we employ the aggregation step from Snorkel framework [35], a widely-used method in the weak supervision community.
Ultimately, the ensemble filtering decisions yield a curated subset that can be reused in later training runs. Moreover, we can use their statistics to estimate training dataset quality by either analyzing retention rate (fraction of samples retained) or averaged mimic score. We discuss weak supervision setups and summarize algorithms in both online and offline stages in Appendix B.